How much is the power of an appliances of 500hm resistance working with 200v supply


1
Expert's answer
2021-04-23T11:06:56-0400

Explanations & Calculations


  • The power of a device is basically given by the equation "\small P=Vi".
  • And with the use of the equation "\small V=iR", another two ways of expressing the same equation can be obtained as "P=\frac{V^2}{R}\qquad P=i^2 R"
  • According to the available data, any suitable form should be chosen.


  • Knowns — "\small V= 200V" and "\small R= 50\,\Omega"


  • Then using the form: "\small P=\frac{V^2 }{R}", the power could be found.

"\qquad\qquad
\begin{aligned}
\small P &=\small \frac{(200V)^2}{50\Omega}\\
&=\small \bold{800\,W}
\end{aligned}"